OpenStudy (anonymous):

directrix =-2 and focus is at (4,3) find equation and the vertex

OpenStudy (anonymous):

@freckles

OpenStudy (jdoe0001):

x=-2?

OpenStudy (anonymous):

vertex = -2

OpenStudy (jdoe0001):

-2? so..hmmm shouldn't it be an ordered pair? 0,-2? -2,0?

OpenStudy (anonymous):

actually i think the directrix is -2

OpenStudy (jdoe0001):

well then, you found it already then =)

OpenStudy (anonymous):

i need help finding the eq and vertex

OpenStudy (jdoe0001):

well, the directrix should have an equation x = -2? y = -2?

OpenStudy (anonymous):

x = -2

OpenStudy (jdoe0001):

 so.. where do you think the vertex would be?

OpenStudy (anonymous):

1,3

OpenStudy (anonymous):

how do i find equation

OpenStudy (jdoe0001):

well, right.. 1,3 is correct.. the equation is ahemm notice the distance from the vertex to either, the focus or directrix, is 3 units thus p = 3 and it's also a horizontal parabola thus \(\bf \begin{array}{llll} (y-{\color{blue}{ k}})^2=4{\color{purple}{ p}}(x-{\color{brown}{ h}}) \\ (x-{\color{brown}{ h}})^2=4{\color{purple}{ p}}(y-{\color{blue}{ k}})\\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ h}},{\color{blue}{ k}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array}\qquad thus \\ \quad \\ \begin{array}{llll} (y-{\color{blue}{ 3}})^2=4{\color{purple}{ (3)}}(x-{\color{brown}{ 1}}) \\ \end{array} \qquad \begin{array}{llll} vertex\ ({\color{brown}{ 1}},{\color{blue}{ 3}})\\ {\color{purple}{ p}}=\textit{distance from vertex to }\\ \qquad \textit{ focus or directrix} \end{array}\)

OpenStudy (jdoe0001):

because is a horizontal parabola, the "y" is the squared variable

OpenStudy (jdoe0001):

would... but need to dash in 1 sec :|

OpenStudy (anonymous):

okay